Using (x + a) (x + b) = x^2 + (a + b) x + ab, find
103 x 98

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

103 x 98
= (100 + 3)(100 - 2)
= 1002+ (3-2)100 - 6
= 10000 + 100 - 6
= 10094

Thanks